# Aliasing and Label-Independent Decomposition of Risk: Beyond the bias–variance trade-off

### v1

@@ II.1.2 Aliasing and Invertibility Error
Consider the situation where a given decomposition $\Theta=\mathcal{M}\oplus\mathcal{U}$ of parameter space with design matrix $\mathbf{M}_{\mathcal{T}\mathcal{M}}$ and nescience $\mathbf{M}_{\mathcal{T}\mathcal{U}}$ is changed by moving one column $\bm{\varphi}$ out of the nescience matrix and into the design matrix. This always decreases (or rather never increases) the norm $\|\mathbf{M}_{\mathcal{T}\mathcal{U}}\|$ but the effect on $\|\mathbf{M}_{\mathcal{T}\mathcal{M}}^{+}\|$ is determined primarily by whether $\bm{\varphi}$ is linearly independent from the other columns of $\mathbf{M}_{\mathcal{T}\mathcal{M}}$ or not, as described in the following theorem (proved in Section IV.IV.2).
@@
II.1.4 Analysis of $\|\mathbf{E}_{\mathbf{B}}\|$ and $\|\mathbf{E}_{\mathbf{B}}\,\bm{\theta}\|$
It is straightforward (see Section IV.IV.3 for details) to show that $\mathbf{E}_{\mathbf{B}}$ can be written as
$\mathbf{E}_{\mathbf{B}}=\begin{pmatrix}\mathbf{P}_{\mathcal{N}}&0\\ 0&\mathbf{I}_{\mathcal{U}}\end{pmatrix},$ (36)
where $\mathbf{P}_{\mathcal{N}}$ is the orthogonal projection of $\mathcal{M}$ onto the null space $\mathcal{N}\subseteq\mathcal{M}$ of $\mathbf{M}_{\mathcal{T}\mathcal{M}}$, and where $\mathbf{I}_{\mathcal{U}}$ is the identity operator on the nescient space $\mathcal{U}$. This implies that its induced operator norm is always [^5] Except in the rare and irrelevant case that $\dim(\mathcal{U})=0$ equal to $1$, and the norm of the product $\mathbf{E}_{\mathbf{B}}\bm{\theta}$ is bounded by the norm of $\bm{\theta}$:
$\|\mathbf{E}_{\mathbf{B}}\bm{\theta}\|\leq\|\mathbf{E}_{\mathbf{B}}\|\|\bm{\theta}\|=\|\bm{\theta}\|.$ (37)
To make more fine-grained statements about the dependence of $\mathbf{E}_{\mathbf{B}}\bm{\theta}$ on the model dimension $m$, we write
where we have made the $m$-dependence explicit. Notice that the first term $\|\mathbf{P}_{\mathcal{N}}\bm{\theta}_{\mathcal{M}}\|^{2}=\|\bm{\theta}_{\mathcal{M}}-\hat{\bm{\theta}}_{\mathcal{M}}\|^{2}$ is the (square of the) bias in $\bm{\theta}_{\mathcal{M}}$ while the second term is exclusively from the nescient parameters, $\bm{\theta}_{\mathcal{U}}$. Since $\mathcal{N}(m)\subseteq\mathcal{N}(m+1)$, it follows that the contribution $\|\mathbf{P}_{\mathcal{N}}\bm{\theta}_{\mathcal{M}}\|$ from parameter bias is a nondecreasing function of $m$. In contrast, the contributions from the nescient parameters is nonincreasing, since it decreases by exactly $|\theta_{m+1}|^{2}$ at each step. Depending on the structure of the parameter vector, the full invertibility error may have nonmonotonic behavior due to the interplay between these contributions.
We first introduce the unstructured model ansatz in which components of $\bm{\theta}$ are independent, identically distributed random variables with mean zero and variance $\sigma^{2}$. [^6] The unstructured model is sensible only when the parameter space has finite dimension, otherwise the norm $\|\bm{\theta}\|$ would be infinite. In this setting the expected invertibility error is
$\mathbb{E}_{\theta}\left[\|\mathbf{E}_{\mathbf{B}}\bm{\theta}\|^{2}\right]$ $=\sigma^{2}\left(\operatorname{Tr}\mathbf{P}_{\mathcal{N}}+\operatorname{Tr}\mathbf{I}_{\mathcal{U}}\right)$ (39)
@@
At each step, $\dim\mathcal{U}$ decreases by one, while $\dim\mathcal{N}$ increases by either zero or one so the total invertibility error is a strictly nonincreasing function of $m$.
In the generic case, $\dim\mathcal{N}(m)=0$ when the number $n$ of rows (data points) is less than the number $m$ of columns (model parameters). This is the well-known fact that the ordinary least squares estimator is unbiased. However, for $m>n$, $\dim\mathcal{N}(m+1)=\dim\mathcal{N}(m)+1$, so parameter bias grows monotonically beyond the interpolation threshold and, on average, cancels the decrease from the nescient parameters.
In contrast to the unstructured case, modelers often have prior information about which parameters are most important and preferentially order the parameter vector to reflect this. In such cases, the total parameter error $\|(\mathbf{E}_{\mathbf{A}}+\mathbf{E}_{\mathbf{B}})\bm{\theta}\|$ tends to be dominated by the invertibility error $\|\mathbf{E}_{\mathbf{B}}\bm{\theta}\|$ for very small and very large model size $m$, where aliasing $\|\mathbf{E}_{\mathbf{A}}\bm{\theta}\|$ is small. For very small $m$, as $m$ increases there is often an initial descent of parameter error due to fewer nescient terms and the model’s increasing ability to capture the signal faithfully. This is conceptually analogous to reducing bias in the classical bias–variance paradigm. However, for very large models, the parameter bias can grow faster than the decrease in the nescient contribution, resulting in growing $\|\mathbf{E}_{\mathbf{B}}\bm{\theta}\|$ as depicted in Fig. 1. This growing error for large models is not analogous to variance and cannot be termed over-fitting. Rather, it reflects the lack of invertibility for large models, specifically, larger parameter bias as more of the mass of $\bm{\theta}$ is projected into the kernel $\mathcal{N}$ of the design matrix $\mathbf{M}_{\mathcal{T}\mathcal{M}}$.
We refer to the phenomenon of nonmonotonic behavior in the invertibility error as over-modeling. Over-modeling is the dominant error in the asymptotic limit of $m\rightarrow\infty$, as contributions from both aliasing and nescience vanish almost surely. The over-modeling phenomenon has largely been overlooked in the double descent literature because it is absent in the unstructured model approach used by most theoretical studies. The example in subsection II.4 below illustrates that unstructured models tend to have optimal performance in the asymptotic limit, while models that exploit prior information often have optimal performance at the classical “sweet spot,” due to asymptotic over-modeling.
II.2 Some Examples
Although the motivating example in Section I.1 for the aliasing decomposition was focused on one-dimensional polynomials, this decomposition applies much more generally to the problem of fitting a function $f:\mathbb{R}^{d}\to\mathbb{R}$ or $f:\mathbb{R}^{d}\to\mathbb{C}$ for arbitrary $d$. We illustrate this with examples of two different choices of models applied to three different data sets. The two bases are random Fourier features (RFF) and random ReLU features (RRF) (described in [9] and [32]).
@@ II.2 Some Examples
II.3 Why ‘aliasing’? Discrete Fourier series
@@
$\begin{pmatrix}\cdots&\mathbf{I}_{n}&\mathbf{I}_{n}&\mathbf{I}_{n}&\cdots\end{pmatrix};$ (45)
@@
II.4 Real-World Example: Cluster Expansion
@@
We first consider the case where there is no natural ordering to either the data or parameters by randomizing rows and columns of $\mathbf{M}$, in line with the unstructured model ansatz above. The resulting behavior (left panel of Fig. 4) aligns with most presentations of double descent in the literature. Before the interpolation threshold, the behavior of the risk curve is the expected U-shape of the classical bias/variance trade-off. Beyond the interpolation threshold, the norms drop, consistent with the generalized aliasing analysis. Furthermore, the optimal model is not at the classical sweet spot but in the asymptotic modern regime since there is little over-modeling.
@@ III Discussion
III.1 General Insights into Modeling
III.1.1 Choosing the Basis
@@
In the discrete Fourier series example (Section II.II.3), the norm of $\mathbf{A}$ is always $1$ and does not decrease to 00 because the columns of $\mathcal{M}$ are specially tuned to the training set to make $\mathbf{M}_{\mathcal{T}\mathcal{U}}$ consist of infinitely many copies of $\mathbf{M}_{\mathcal{T}\mathcal{M}}$. This aligning of the basis functions to sample points explains why extreme over-fitting is rarely a problem in discrete Fourier transforms, in spite of it being formally equivalent to ordinary least squares regression at the interpolation threshold.
III.1.2 Choosing Training Points
@@ III.2 Regularization
where $n$ is the the number of training points and $\lambda$ is a user-chosen parameter. As shown in Section IV.IV.4, this is equivalent to the alternative problem of minimizing
@@ III.3 Outlook
An important open question is: Under what conditions is the asymptotic risk less than that of the classical “sweet spot”? The preceding analysis has sharpened that question to: When will there be over-modeling? Random feature models, such as in Figure 2, but presumably also neural networks and other machine learning models, are approximately unstructured, so they do not exhibit over-modeling and are generically most effective in the over-parameterized, modern regime. In contrast, physics-based models are most effective in the classical regime, where they leverage prior knowledge.
@@ IV Methods
IV.1 Interleaving of Eigenvalues in Rank-one Updates
@@
Theorem IV.1 also leads us to the following fundamental result for analyzing the operator norms of the aliasing and invertibility operators $\mathbf{A}$ and $\mathbf{B}$ at least in the finite dimensional case.
Theorem IV.2.
@@
Expanding $\widetilde{\mathbf{X}}\widetilde{\mathbf{X}}^{\mathsf{T}}$ gives $\widetilde{\mathbf{X}}\widetilde{\mathbf{X}}^{\mathsf{T}}=\mathbf{X}\mathbf{X}^{\mathsf{T}}+\bm{\varphi}\bm{\varphi}^{\mathsf{T}}$, where $\bm{\varphi}\bm{\varphi}^{\mathsf{T}}$ is positive semidefinite, so Theorem IV.1 implies that
@@ IV.2 Norm of $\mathbf{A}$
IV.2.1 Nescience:
First consider what happens to the nescience matrix $\mathbf{M}_{\mathcal{T}\mathcal{U}}$ when a column $\bm{\varphi}_{m+1}$ is removed from $\mathbf{Y}(m)=[\bm{\varphi}_{m}\quad\mathbf{Y}(m+1)]$. Expanding the product $\mathbf{Y}(m)\mathbf{Y}(m)^{\mathsf{T}}$ gives $\mathbf{Y}(m)\mathbf{Y}(m)^{\mathsf{T}}=\bm{\varphi}_{m+1}\bm{\varphi}_{m+1}^{\mathsf{T}}+\mathbf{Y}(m+1)\mathbf{Y}(m+1)^{\mathsf{T}}$. Since $\bm{\varphi}_{m+1}\bm{\varphi}_{m+1}^{\mathsf{T}}$ is positive semidefinite, Theorem IV.1 applies and guarantees that the norms satisfy $\|\mathbf{Y}(m)\|\geq\|\mathbf{Y}(m+1)\|$, and thus the norm $\|\mathbf{Y}(m)\|$ is a nonincreasing function of $m$.
IV.2.2 Pseudoinverse of Design:
Consider now the pseudoinverse term $\|\mathbf{M}_{\mathcal{T}\mathcal{M}}^{+}\|$ when $\bm{\varphi}_{m+1}$ is adjoind to $\mathbf{X}(m)$ to create $\mathbf{X}(m+1)$. Theorem IV.2 guarantees that whenever $\bm{\varphi}_{m+1}$ is linearly independent of the old model (does not lie in the column space of $\mathbf{X}(m)$), then the induced norm of the new pseudoinverse is bounded below by the induced norm of the old pseudoinverse:
@@ IV.2.2 Pseudoinverse of Design:
IV.2.3 Limiting behavior of $\mathbf{A}$
@@
IV.3 Norm of $\mathbf{E}_{\mathbf{B}}$ and $\mathbf{E}_{\mathbf{B}}\bm{\theta}$
@@
The matrix $V_{1}$ consists of the first $r$ columns of an orthonormal matrix $V$, i.e., $V=[V_{1}|V_{2}]$, with $VV^{\mathsf{T}}=V_{1}V_{1}^{\mathsf{T}}+V_{2}V_{2}^{\mathsf{T}}=\mathbf{I}$. Hence Theorem IV.1 guarantees that
@@
As discussed in Section II.II.1.4, to better understand the dependence of $\mathbf{E}_{\mathbf{B}}\bm{\theta}$ on the model dimension $m$, we decompose the square of the norm $\|\mathbf{E}_{\mathbf{B}}\|^{2}$ further into model invertibility and nescience contributions

### v2

@@ II.3.3 Generalized Aliasing
Recall from Eq. (24) that aliasing operator is the product of the inverse design matrix $\mathbf{M}_{\mathcal{T}\mathcal{M}}^{+}$ and the nescient transformation $\mathbf{M}_{\mathcal{T}\mathcal{U}}$. Increasing the number of model parameters by moving one column $\bm{\varphi}$ out of the nescience matrix and into the design matrix always decreases (or rather never increases) the norm $\|\mathbf{M}_{\mathcal{T}\mathcal{U}}\|$. However, the effect on $\|\mathbf{M}_{\mathcal{T}\mathcal{M}}^{+}\|$ is determined primarily by whether $\bm{\varphi}$ is linearly independent from the other columns of $\mathbf{M}_{\mathcal{T}\mathcal{M}}$ or not, as described in the following theorem (proved in Section V.2).
@@
II.3.4 Parameter-Data Trade-off and Over-Modeling
Because the data and parameter insufficiency terms are non-decreasing and non-increasing respectively, there is an inherent trade-off between them. To study this trade-off, we introduce the combined parameter and data insufficiency error:
$\ \|\mathbf{E}_{I}\bm{\theta}\|^{2}$ $=\|\mathbf{P}_{\mathcal{N}(m)}\bm{\theta}_{\mathcal{M}(m)}\|^{2}+\|\bm{\theta}_{\mathcal{U}(m)}\|^{2}.$ (31)
where we have made the $m$-dependence explicit. To make statements about the dependence of the combined insufficiency errors, we consider two prior distributions for the distribution of the components of $\bm{\theta}$.
We first introduce the random feature model ansatz in which components of $\bm{\theta}$ are independent, identically distributed random variables with mean zero and variance $\sigma^{2}$. [^6] The random feature model is sensible only when the parameter space has finite dimension, otherwise the norm $\|\bm{\theta}\|$ would be infinite. In this setting the total expected insufficiency error is
$\mathbb{E}_{\theta}\left[\|\mathbf{E}_{I}\bm{\theta}\|^{2}\right]$ $=\sigma^{2}\left(\operatorname{Tr}\mathbf{P}_{\mathcal{N}}+\operatorname{Tr}\mathbf{I}_{\mathcal{U}}\right)$ (32)
@@ II.3.4 Parameter-Data Trade-off and Over-Modeling
At each step, $\dim\mathcal{U}$ decreases by one, while $\dim\mathcal{N}$ increases by either zero or one so the total insufficiency error is a strictly nonincreasing function of $m$.
In contrast to the random feature case, modelers often have prior information about which parameters are most important and preferentially order the parameter vector to reflect this. In such cases and for very small $m$, as $m$ increases there is often an initial descent of parameter error due to fewer nescient terms and the model’s rapidly increasing ability to capture the signal faithfully. This is conceptually analogous to reducing bias in the classical bias–variance paradigm. However, for very large models, the data insufficiency grows faster than the decrease in the nescient contribution. This growing error for large models is not analogous to variance and cannot be termed over-fitting. Rather, it reflects the lack of invertibility for large models, specifically, larger parameter bias as more of the mass of $\bm{\theta}$ is projected into the kernel $\mathcal{N}$ of the design matrix $\mathbf{M}_{\mathcal{T}\mathcal{M}}$.
We refer to the phenomenon of nonmonotonic behavior in the invertibility error as over-modeling. Over-modeling is the dominant error in the asymptotic limit of $m\rightarrow\infty$, as contributions from both aliasing and parameter insufficiency vanish almost surely. The over-modeling phenomenon has largely been overlooked in the double descent literature because it is absent in random feature models used by most theoretical studies. The example in subsection III.3 below illustrates that random feature models tend to have optimal performance in the asymptotic limit, while models that exploit prior information often have optimal performance at the classical “sweet spot,” due to asymptotic over-modeling.
III Demonstrations and Applications
III.1 Random Feature Models
Although the motivating example in Section II.1 for the aliasing decomposition was focused on one-dimensional polynomials, this decomposition applies much more generally to the problem of fitting a function $f:\mathbb{R}^{d}\to\mathbb{R}$ or $f:\mathbb{R}^{d}\to\mathbb{C}$ for arbitrary $d$. We illustrate this with examples of two different choices of models applied to three different data sets. The two bases are random Fourier features (RFF) and random ReLU features (RRF) (described in [8] and [32]).
@@ III.1 Random Feature Models
III.2 Why ‘aliasing’? Discrete Fourier series
@@
$=\begin{pmatrix}\cdots&\mathbf{I}_{n}&\mathbf{I}_{n}&\mathbf{I}_{n}&\cdots\end{pmatrix};$ (38)
@@
III.3 Real-World Example: Cluster Expansion
@@
We first consider the case where there is no natural ordering to either the data or parameters by randomizing rows and columns of $\mathbf{M}$, in line with the random feature model ansatz above. The resulting behavior (left panel of Fig. 5) aligns with most presentations of double descent in the literature. Before the interpolation threshold, the behavior of the risk curve is the expected U-shape of the classical bias/variance trade-off. Beyond the interpolation threshold, the norms drop, consistent with the generalized aliasing analysis. Furthermore, the optimal model is not at the classical sweet spot but in the asymptotic modern regime since there is little over-modeling.
@@ IV Discussion
IV.1 General Insights into Modeling
IV.1.1 Choosing the Basis
@@
In the discrete Fourier series example (Section II.2.III.2), the norm of $\mathbf{A}$ is always $1$ and does not decrease to 00 because the columns of $\mathcal{M}$ are specially tuned to the training set to make $\mathbf{M}_{\mathcal{T}\mathcal{U}}$ consist of infinitely many copies of $\mathbf{M}_{\mathcal{T}\mathcal{M}}$. This aligning of the basis functions to sample points explains why extreme over-fitting is rarely a problem in discrete Fourier transforms, in spite of it being formally equivalent to ordinary least squares regression at the interpolation threshold.
IV.1.2 Choosing Training Points
@@ IV.2 Regularization
where $n$ is the the number of training points and $\lambda$ is a user-chosen parameter. As shown in Section V.V.4, this is equivalent to the alternative problem of minimizing
@@ IV.3 Outlook
An important open question is: Under what conditions is the asymptotic risk less than that of the classical “sweet spot”? The preceding analysis has sharpened that question to: When will there be over-modeling? Random feature models, such as in Figure 3, but presumably also neural networks and other machine learning models, often do not exhibit over-modeling and are generically most effective in the over-parameterized, modern regime. In contrast, physics-based models are most effective in the classical regime, where they leverage prior knowledge.
@@ V Methods
V.1 Interleaving of Eigenvalues in Rank-one Updates
@@
Theorem V.1 also leads us to the following fundamental result for analyzing the operator norms of the aliasing and invertibility operators $\mathbf{A}$ and $\mathbf{B}$ at least in the finite dimensional case.
Theorem V.2.
@@
Expanding $\widetilde{\mathbf{X}}\widetilde{\mathbf{X}}^{\mathsf{T}}$ gives $\widetilde{\mathbf{X}}\widetilde{\mathbf{X}}^{\mathsf{T}}=\mathbf{X}\mathbf{X}^{\mathsf{T}}+\bm{\varphi}\bm{\varphi}^{\mathsf{T}}$, where $\bm{\varphi}\bm{\varphi}^{\mathsf{T}}$ is positive semidefinite, so Theorem V.1 implies that
@@ V.2 Norm of $\mathbf{A}$
V.2.1 Nescience:
First consider what happens to the nescience matrix $\mathbf{M}_{\mathcal{T}\mathcal{U}}$ when a column $\bm{\varphi}_{m+1}$ is removed from $\mathbf{Y}(m)=[\bm{\varphi}_{m}\quad\mathbf{Y}(m+1)]$. Expanding the product $\mathbf{Y}(m)\mathbf{Y}(m)^{\mathsf{T}}$ gives $\mathbf{Y}(m)\mathbf{Y}(m)^{\mathsf{T}}=\bm{\varphi}_{m+1}\bm{\varphi}_{m+1}^{\mathsf{T}}+\mathbf{Y}(m+1)\mathbf{Y}(m+1)^{\mathsf{T}}$. Since $\bm{\varphi}_{m+1}\bm{\varphi}_{m+1}^{\mathsf{T}}$ is positive semidefinite, Theorem V.1 applies and guarantees that the norms satisfy $\|\mathbf{Y}(m)\|\geq\|\mathbf{Y}(m+1)\|$, and thus the norm $\|\mathbf{Y}(m)\|$ is a nonincreasing function of $m$.
V.2.2 Pseudoinverse of Design:
Consider now the pseudoinverse term $\|\mathbf{M}_{\mathcal{T}\mathcal{M}}^{+}\|$ when $\bm{\varphi}_{m+1}$ is adjoind to $\mathbf{X}(m)$ to create $\mathbf{X}(m+1)$. Theorem V.2 guarantees that whenever $\bm{\varphi}_{m+1}$ is linearly independent of the old model (does not lie in the column space of $\mathbf{X}(m)$), then the induced norm of the new pseudoinverse is bounded below by the induced norm of the old pseudoinverse:
@@ V.2.2 Pseudoinverse of Design:
V.2.3 Limiting behavior of $\mathbf{A}$
@@
V.3 Norm of $\mathbf{E}_{\mathbf{B}}$ and $\mathbf{E}_{\mathbf{B}}\bm{\theta}$
@@
The matrix $V_{1}$ consists of the first $r$ columns of an orthonormal matrix $V$, i.e., $V=[V_{1}|V_{2}]$, with $VV^{\mathsf{T}}=V_{1}V_{1}^{\mathsf{T}}+V_{2}V_{2}^{\mathsf{T}}=\mathbf{I}$. Hence Theorem V.1 guarantees that
@@
As discussed in Section II.3, to better understand the dependence of $\mathbf{E}_{\mathbf{B}}\bm{\theta}$ on the model dimension $m$, we decompose the square of the norm $\|\mathbf{E}_{\mathbf{B}}\|^{2}$ further into model invertibility and nescience contributions